–3(10x – 3) + 10(3x – 3) = x – 7
Answer by stanbon(48568)

–3(10x – 3) + 10(3x – 3) = x – 7
Simplify to get:
-30x+9+30x-30=x-7
Combine line terms to get:
-21=x-7
Add 7 to both sides to get:
x=-14
Cheers,
Stan H.